# A Large Language Model Enhanced Sequential Recommender for Joint Video and Comment Recommendation

### v1

@@ 3.1.4. Preference Extraction from Sequence Representations
For the tasks of video and comment recommendation, we employ different methods to extract user interests from sequence representations. Specifically, we exploit additive attention (Bahdanau et al., 2015) in video recommendation to manage the importance of different representations:
(8) $\mathbf{s}^{v}$ $=F_{v}(\widehat{\mathbf{H}}^{v})=\sum_{i=1}^{n}\alpha_{i}\mathbf{h}_{i}^{v},\quad\alpha_{i}=\frac{\text{exp}(f(\mathbf{h}_{i}^{v}))}{\sum_{k=1}^{n}\text{exp}(f(\mathbf{h}_{k}^{v}))},$
@@
(9) $\mathbf{s}^{c}$ $=F_{c}(\widehat{\mathbf{H}}^{c})=\sum_{j=1}^{m}\beta_{j}\mathbf{h}_{j}^{c},\quad\beta_{j}=\frac{\text{exp}(g(\mathbf{e}^{t}_{m+1},\mathbf{h}_{j}^{c}))}{\sum_{k=1}^{m}\text{exp}(g(\mathbf{e}^{t}_{m+1},\mathbf{h}_{k}^{c}))}.$
where $\mathbf{h}_{j}^{c}$ is the $j$-th vector of $\widehat{\mathbf{H}}^{c}$, $g(\cdot)$ is an attention weight function (similar to multi-head attention) with $\mathbf{e}^{t}_{m+1}$ as the query, $\mathbf{s}^{c}$ denotes the representation of user comment preference on video $\tilde{v}_{m+1}$.
@@ 3.2. Enhanced Preference Modeling via LLM
In recommendation scenarios, LLM generally plays two roles: as a text feature encoder (Section 3.1) or a recommender (Zhang et al., 2023b; Bao et al., 2023). The former offers greater flexibility and efficiency, while the latter is beneficial for the more effective utilization of LLM’s excellent semantic understanding and knowledge reasoning capabilities. Therefore, in this section, we further perform enhanced user preference modeling through a LLM recommender to improve the semantics of the SR model. It is worth noting that this component is only included in the alignment training stage and does not exist in the subsequent fine-tuning and practical deployment.
Following previous works (Hou et al., 2023; Dai et al., 2023; Zhang et al., 2023b), we format the user interaction sequences in the form of textual instructions and utilize LLM for generative recommendation. Specifically, the instruction templates for two tasks are as follows:
@@
where $[-1]$ means taking the hidden state corresponding to the final token of $I(\mathcal{S}_{v},\mathcal{S}_{c})$, and $\tilde{\mathbf{s}}$ is the enhanced preference representation from the LLM recommender.
@@ 3.3.1. Sequential-Supplemental Preference Contrast
Input augmentation via comment diversification. In our scenario, there is abundant and diverse comment data, which can be leveraged for data augmentation. Specifically, we apply comment diversification with a certain probability to augment user video and comment interaction histories: (1) For each video interaction, we select two distinct popular comments to pair with the same video as inputs for the SR model and the LLM recommender. (2) For each comment interaction, we sample a couple of virtual interaction comments and mix them with the real interaction comments. Then, the mixed interaction comments are randomly allocated to the inputs of two components.
Preference representation contrastive learning. In the personalized preference alignment stage, we jointly learn video and comment recommendation tasks. For each data instance, the sequential preference representation of the SR model is calculated by Eq. (8) or Eq. (9). For the sake of brevity, here we no longer distinguish between the two tasks, and the representation is denoted as $\mathbf{s}$. Additionally, by the approach mentioned in Section 3.2, we obtain the enhanced preference representation of the LLM recommender, denoted as $\tilde{\mathbf{s}}$. The sequential-supplemental preference contrastive loss can be formulated as:
@@
where $\mathbf{x}$ and $\mathbf{y}^{+}$ denote a pair of positive instances for contrastive learning. $\mathcal{O}_{\mathbf{y}}$ denotes a set of contrastive instances, consisting of both positive and negative samples. $\tau$ is a temperature coefficient, and $\text{cos}(\cdot,\cdot)$ represents the cosine similarity function.
@@ 3.3.2. Video-Comment Preference Contrast
Furthermore, we distinguish between user preferences that are extracted from video and comment sequence representations, based on the enhanced preference representation provided by the LLM recommender. The core idea is that when LLM is instructed to make video recommendations, the preference representation (i.e., $\tilde{\mathbf{s}}$) should be aligned with that extracted from the video sequence (i.e., $\widehat{\mathbf{H}}^{v}$), rather than from the comment sequence (i.e., $\widehat{\mathbf{H}}^{c}$), and vice versa. Formally, we denote the instruction/task objective (video or comment recommendation) of the instance as $o\in\{v,c\}$. The video-comment preference contrastive loss is formulated as follows:
@@ 3.3.3. Alignment Training Objective
where $\lambda$ and $\mu$ are hyperparameters for the trade-off between various objectives.
@@ 3.4. Recommendation-Oriented Fine-tuning
After the personalized preference alignment stage, we further fine-tune the alignment-enhanced SR model for video and comment recommendation tasks. The supplemental LLM recommender will be discarded during the fine-tuning stage and practical deployment.
@@ 3.4.1. Fine-tuning
(18) $\mathcal{L}_{VRT}$ $=\operatorname{InfoNCE}(\mathbf{s}^{v},(\mathbf{e}^{id}_{n+1}+\mathbf{e}^{t}_{n+1}),\mathcal{V}),$
where $\mathcal{V}$ is the entire set of candidate videos. $\mathbf{e}^{t}_{n+1}=\operatorname{LLM}(t_{n+1})$ represents the title embedding of the next video $v_{n+1}$.
Finally, our overall optimization loss function in the fine-tuning stage can be denoted as follows:
(19) $\mathcal{L}_{FT}=\mathcal{L}_{VRT}+\eta\mathcal{L}_{REG},$
where $\eta$ is the hyperparameter for the regularization loss. Regarding the comment recommendation task, we use comment preference representation $\mathbf{s}^{c}$ for prediction, and all candidates (i.e., $\mathcal{C}_{\tilde{v}_{m+1}}$) are selected from comments of video $\tilde{v}_{m+1}$ that the user is watching.
@@ 3.4.2. Time Complexity Analysis
In practical deployment, we retain the fine-tuned SR model for recommendation, and discard the supplemental LLM recommender. Besides, the textual information associated with the videos and comments is encoded offline to be directly accessed during online inference. Therefore, we just discuss the time complexity of the subsequent sequential user preference modeling process. Formally, the time consumption of Transformer layers in Eq. (4) and Eq. (5) is $O(n^{2}d+m^{2}d)$, where $d$ denotes the model dimension, $n$ and $m$ denote the lengths of user video and comment interaction histories, respectively. Furthermore, $O(nd^{2}+md^{2})$ is the complexity of intermediate linear operations. And the time used by fusion encoders in Eq. (6) and Eq. (7) is $O(mnd)$. In addition, some low-order time complexities are ignored, such as the $\mathcal{O}(nd)$ or $\mathcal{O}(md)$ in preference extraction (Section 3.1.4). Overall, the time complexity of sequential user preference modeling is $\mathcal{O}(n^{2}d+m^{2}d+nd^{2}+md^{2}+mnd)$. In the end, the time consumed for calculating scores with candidates is $\mathcal{O}(Nd)$, where $N$ denotes the number of candidates. In general, when it comes to online deployment, our LSVCR model has comparable time complexity to mainstream sequential recommendation models (e.g., SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019)).
@@ 4. Experiment
In this section, we first set up the experiments and then present the results and analyses of our proposed LSVCR.
@@ 4.1.1. Dataset
We evaluate the proposed framework on the large-scale industrial dataset constructed from a popular Chinese micro-video platform KuaiShou for video and comment recommendation tasks. To be specific, the dataset is built from the interaction logs of 19,691 users from October 24 to October 31, 2023. For data preprocessing, we split the user interactions into training, validation, and testing sets according to the timestamp. The data from the last two days is used for validation and testing respectively, while the remaining data is utilized to train our framework. The detailed statistics of the dataset are summarized in Table 1.
@@ 4.1.2. Baseline Models
[…]
Comment recommendation. The baseline methods for comment recommendation include: (1) Traditional recommendation model: DSSM (Huang et al., 2013) uses dual encoders to separately encode user and candidate comments. (2) Sequential modeling methods: GRU (Cho et al., 2014) intuitively utilizes GRU to derive short- and long-term user representations (An et al., 2019; Hidasi et al., 2016; Li et al., 2017). ATT (Vaswani et al., 2017) employs an additive attention (ATT) mechanism to model the user comment interaction sequence (Wu et al., 2019b). MHA (Vaswani et al., 2017) adopts multi-head attention (MHA) and additive attention to obtain the representation of the user comment interaction history (Wu et al., 2019b, c). (3) Query-based models: ZAM (Ai et al., 2019) introduces a zero vector into the attention mechanism to dynamically control the influence of user interaction history on product search. TEM (Bi et al., 2020) uses Transformer to encode the query and user interaction sequence for more flexible personalization adaptation. For all comment recommendation baselines, we adopt trainable BERT (Devlin et al., 2019) as their text encoder. Furthermore, we consider the title embedding of the video being watched by the user as the query for query-based models.
Table 2. Performance comparison of different methods for video recommendation. The best and second-best results are highlighted in bold and underlined font, respectively.
@@
[…]
UniSRec 0.2570 0.3159 0.1875 0.2135 0.1737
LSVCR 0.2719 0.3322 0.2037 0.2233 0.1893
Table 3. Performance comparison of various models on the comment recommendation task.
@@ 4.1.3. Evaluation Settings
To evaluate the performance of video and comment recommendation tasks, we adopt three widely used metrics Recall@$K$, NDCG@$K$, and MRR, where $K$ is set to 5 and 10. We perform negative sampling evaluations for both tasks. For video recommendation, we pair the ground-truth item with 99 randomly sampled items. Regarding comment recommendation, the candidate number is limited to 100. If the number of comments below the current video is less than 100, all comments will be used. Conversely, negative sampling will be conducted until there are 100 candidates.
@@ 4.1.4. Implementation Details
For the personalized preference alignment stage, we utilize ChatGLM3 (Du et al., 2022; Zeng et al., 2023) as our LLM backbone and perform low-rank adaptation based on LoRA (Hu et al., 2022). We employ the AdamW optimizer with a learning rate of 3e-4 for optimization. With the application of data parallelism and gradient accumulation, the total batch size reaches 128. The loss coefficients, $\lambda$ and $\mu$, are set to 1 and 0.5 respectively. The temperature coefficient $\tau$ is assigned a value of 0.07. The maximum lengths of user video and comment interaction histories are set to 20 for the alignment stage and 50 for the fine-tuning stage. Finally, we conduct a total of 5000 alignment steps. For the recommendation-oriented fine-tuning stage, the embedding dimension is set to 64, and the learning rate is tuned in {0.005, 0.003, 0.001, 0.0005, 0.0001}, which are consistent with all baseline models. The hyperparameter $\eta$ is tuned in {0.1, 0.5 }. More details can be found in our implementation code [^2] https://github.com/RUCAIBox/LSVCR/.
Table 4. Ablation study of LSVCR on the tasks of video and comment recommendation.

### v2

@@ 4.1.4. Preference Extraction from Sequence Representations
For the tasks of video and comment recommendation, we employ different methods to extract user preferences from sequence representations. Specifically, we exploit additive attention (Bahdanau et al., 2015) in video recommendation to manage the importance of different representations:
(7) $\mathbf{s}^{v}$ $=F_{v}(\widehat{\mathbf{H}}^{v})=\sum_{i=1}^{n}\alpha_{i}\mathbf{h}_{i}^{v},\ \ \ \ \alpha_{i}$ $=\frac{\text{exp}(f(\mathbf{h}_{i}^{v}))}{\sum_{k=1}^{n}\text{exp}(f(\mathbf{h}_{k}^{v}))},$
@@
(8) $\mathbf{s}^{c}$ $=F_{c}(\widehat{\mathbf{H}}^{c})=\sum_{j=1}^{m}\beta_{j}\mathbf{h}_{j}^{c},\ \ \ \ \beta_{j}$ $=\frac{\text{exp}(g(\mathbf{e}^{t}_{m+1},\mathbf{h}_{j}^{c}))}{\sum_{k=1}^{m}\text{exp}(g(\mathbf{e}^{t}_{m+1},\mathbf{h}_{k}^{c}))}.$
where $g(\cdot,\cdot)$ is an attention weight function (similar to the attention function in Transformer) with $\mathbf{e}^{t}_{m+1}$ as the query, $\mathbf{s}^{c}$ denotes the representation of user comment preference on video $\tilde{v}_{m+1}$.
@@ 4.2. Enhanced Preference Modeling via LLM
In recommendation scenarios, LLM generally plays two roles: as a text feature encoder (Section 4.1) or a recommender (Zhang et al., 2023b; Bao et al., 2023). The former offers greater flexibility and efficiency, while the latter is beneficial for maximizing LLM’s excellent semantic understanding and knowledge reasoning capabilities. Therefore, we further introduce enhanced user preference modeling through a LLM recommender to improve the semantics of the SR model. Notably, this component is only included in the alignment training stage and does not exist in the subsequent fine-tuning and practical deployment.
Following previous works (Hou et al., 2023b; Dai et al., 2023; Zhang et al., 2023b), we format the user interaction behaviors in the form of textual instructions and utilize LLM for generative recommendation. Specifically, the instruction templates for two tasks are as follows:
{adjustwidth}
-7mm-7mm
@@
where $[-1]$ means taking the hidden state of the final token, $\tilde{\mathbf{s}}$ is the enhanced preference representation from the LLM recommender.
@@ 4.3.1. Sequential-Supplemental Preference Contrast
Input augmentation via comment diversification. In our context, apart from the comments that users have interacted with, there is also a vast amount of comment data. To improve data utilization, we introduce comment diversification for data augmentation. Specifically, we apply hot comments to augment user video and comment interaction histories with a certain probability: (1) For each video interaction record, we select two distinct popular comments to pair with the same video as inputs for the SR model and the LLM recommender. (2) For a given comment interaction record, assume that it contains k comments that the user has interacted with. We first sample k popular comments under the video as virtual interaction comments and mix them with the actual interaction comments. After that, we randomly allocate mixed interaction comments to two groups, each containing k comments. These two groups will be used to construct inputs for the SR model and the LLM recommender respectively.
Preference representation contrastive learning. In the personalized preference alignment stage, we jointly learn video and comment recommendation tasks. For each data instance, the sequential preference representation of the SR model is calculated by Eq. (7) or Eq. (8) according to the task objective. To simplify, here we no longer distinguish between the two tasks, and the representations are uniformly denoted as $\mathbf{s}$. Additionally, by the approach mentioned in Section 4.2, we obtain the enhanced preference representation of the LLM recommender (different task instances use different instruction templates), denoted as $\tilde{\mathbf{s}}$. The sequential-supplemental preference contrastive loss can be formulated as:
@@
where $\mathbf{x}$ and $\mathbf{y}^{+}$ denote a pair of positive instances for contrastive learning. $\mathcal{O}_{\mathbf{y}}$ denotes contrastive instances consisting of both positive and negative samples. $\tau$ is a temperature coefficient.
@@ 4.3.2. Video-Comment Preference Contrast
Furthermore, we disentangle user preferences derived from video and comment interaction sequences, based on the enhanced preference representation provided by the LLM recommender. The core idea is that when LLM is instructed to make video recommendations, the preference representation (i.e., $\tilde{\mathbf{s}}$) should be aligned with that extracted from the video sequence (i.e., $\widehat{\mathbf{H}}^{v}$), rather than from the comment sequence (i.e., $\widehat{\mathbf{H}}^{c}$), and vice versa. Formally, we denote the task objective (video or comment recommendation) of an instance as $o\in\{v,c\}$. The video-comment preference contrastive loss is as follows:
@@ 4.3.3. Alignment Training Objective
where $\lambda$ and $\mu$ are hyperparameters that balance various objectives.
@@ 4.4. Recommendation-Oriented Fine-tuning
After the personalized preference alignment stage, we further fine-tune the alignment-enhanced SR model for video and comment recommendation tasks respectively. The supplemental LLM recommender will be discarded during the fine-tuning and deployment.
@@ 4.4.1. Fine-tuning
(17) $\mathcal{L}_{VR}$ $=\operatorname{InfoNCE}(\mathbf{s}^{v},(\mathbf{e}^{id}_{n+1}+\mathbf{e}^{t}_{n+1}),\mathcal{V}),$
where $\mathcal{V}$ is the entire set of candidate videos. $\mathbf{e}^{t}_{n+1}=\operatorname{LLM}(t_{n+1})$ represents the title embedding of the next video $v_{n+1}$. Finally, our overall optimization loss function for video recommendation in the fine-tuning stage can be denoted as follows:
(18) $\mathcal{L}_{VFT}=\mathcal{L}_{VR}+\eta\mathcal{L}_{REG},$
where $\eta$ is the hyperparameter for the regularization loss.
Regarding the comment recommendation task, we employ $\mathbf{s}^{c}$ (Eq. (8)) for prediction, and all candidates (i.e., $\mathcal{C}_{\tilde{v}_{m+1}}$) are selected from comments of video $\tilde{v}_{m+1}$ that the user is watching. Consequently, the objective of comment recommendation is as follows:
(19) $\mathcal{L}_{CR}=\operatorname{InfoNCE}(\mathbf{s}^{c},\mathbf{e}^{c}_{m+1},\mathcal{C}_{\tilde{v}_{m+1}}),$
where $\mathbf{e}^{c}_{m+1}=\operatorname{LLM}(c_{m+1})$ denotes the positive comment embedding that encoded by LLM. After combining with the regularization loss, we can also obtain the loss for comment recommendation during fine-tuning, denoted as $\mathcal{L}_{VFT}$.
@@ 4.4.2. Time Complexity Analysis
In practical deployment, we retain the fine-tuned SR model for recommendation, and discard the supplemental LLM recommender. Besides, the textual information associated with the videos and comments is encoded offline to be directly accessed during online inference. Therefore, we just discuss the time complexity of the subsequent sequential user preference modeling process. In particular, the time consumption of the SR model primarily involves encoding the two interaction sequences of video and comments (Eq. (4)). This results in a time consumption of $O(ln^{2}d+lm^{2}d)$, where $l$ denotes the number of Transformer layers, $d$ denotes the model dimension, $n$ and $m$ denote the lengths of user video and comment interaction histories, respectively. In contrast, the time complexity of directly using LLM as a recommender is $O(LN^{2}D)$, where $L$ is the layer number, $H$ denotes the hidden state size, and $N$ is the length of the input text. Since $L>>l$ and $H>>d$, this makes the deployment of the LLM recommender far more resource-intensive than our approach. Compared to the text-based method FSDA (Zhang et al., 2019) with a time complexity of $O(2ln^{2}d)$ and the multi-behavior method DMT (Gu et al., 2020) with a time consumption of $O(ln^{2}d+lm^{2}d)$, our approach has a comparable time complexity.
@@ 5.1.1. Dataset
We evaluate the proposed framework on the large-scale industrial dataset constructed from KuaiShou platform for video and comment recommendation tasks [^1] https://github.com/lyingCS/KuaiComt.github.io (Zhang et al., 2025). To be specific, the dataset is built from the interaction logs of 19,691 users from October 24 to October 31, 2023. For data preprocessing, we split the user interactions into training, validation, and testing sets according to the timestamp. The data from the last two days is used for validation and testing respectively, while the remaining data is utilized to train our framework. The detailed statistics of the dataset are summarized in Table 1.
@@ 5.1.2. Baseline Models
Video recommendation. The baselines for video recommendation can be divided into three categories: (1) Traditional sequential recommendation models: Caser (Tang and Wang, 2018) employs CNN to model user preference through horizontal and vertical convolutional filters. GRU4Rec (Hidasi et al., 2016) is a RNN-based model that utilizes GRU to encode historical behavior sequences. SASRec (Kang and McAuley, 2018) uses a unidirectional Transformer for next item prediction. BERT4Rec (Sun et al., 2019) adopts a bidirectional Transformer to model the item sequence. NARM (Li et al., 2017) combines the GRU with an attention mechanism to capture users’ interests. (2) Multi-behavior models: DMT (Gu et al., 2020) exploits multiple Transformers with MoEs to model diverse user behavior sequences. MBHT (Yang et al., 2022) empowers the Transformer with a multi-behavior hypergraph to capture diverse multi-behavior dependencies. (3) Text-enhanced models: FDSA (Zhang et al., 2019) applies separate self-attention networks to model item-level and feature-level sequences. S3-Rec (Zhou et al., 2020) proposes pre-training with mutual information maximization to learn the correlation between items and attributes. UniSRec (Hou et al., 2022) incorporates textual item embeddings with a MoE adapter to learn universal representations. We use both item ID and text. VQ-Rec (Hou et al., 2023a) learns vector-quantized representations for sequential recommendation.
Comment recommendation. The baselines for comment recommendation include: (1) Traditional recommendation model: DSSM (Huang et al., 2013) uses dual encoders to separately encode user and candidate comments. (2) Sequential modeling methods: GRU intuitively utilizes GRU (Cho et al., 2014) to derive short- and long-term user representations (An et al., 2019; Hidasi et al., 2016; Li et al., 2017). ATT employs an additive attention (Vaswani et al., 2017) mechanism to model the user comment interaction sequence (Wu et al., 2019b). MHA adopts multi-head attention (Vaswani et al., 2017) and additive attention to obtain the representation of the user comment interaction history (Wu et al., 2019b, c). UniSRecC (Hou et al., 2022) mirrors the model structure of UniSRec for comment modeling (3) Query-based models: ZAM (Ai et al., 2019) introduces a zero vector into the attention mechanism to dynamically control the influence of user interaction history on product search. TEM (Bi et al., 2020) uses Transformer to encode the query and user interaction sequence for more flexible personalization adaptation. For all comment recommendation baselines, we adopt trainable BERT (Devlin et al., 2019) as their text encoder. Furthermore, we consider the title embedding of the video being watched by the user as the query for query-based models.
Table 2. Performance comparison of video recommendation. The best and second-best results are highlighted in bold and underlined font, respectively. “*” denotes that the improvements are significant at the level of 0.01 with paired $t$-test.
@@
S3-Rec 0.2226 0.2899 0.1619 0.1837 0.1509
UniSRect+ID 0.2570 0.3159 0.1875 0.2135 0.1737
VQ-Rec 0.2540 0.3182 0.1813 0.2150 0.1759
LSVCR 0.2719* 0.3322* 0.2037* 0.2233* 0.1893*
Table 3. Performance comparison on the industrial dataset for comment recommendation. “*” denotes that the improvements are significant at the level of 0.01 with paired $t$-test.
@@ 5.1.3. Evaluation Settings
To evaluate the performance of video and comment recommendation tasks, we adopt three widely used metrics Recall@$K$, NDCG@$K$, and MRR, where $K$ is set to 5 and 10. For a fair comparison, the maximum historical sequence length is uniformly set to 50 for all comparison methods. For video recommendation, we pair the ground-truth item with 99 randomly sampled items. Regarding comment recommendation, the candidate number is limited to 100. If the number of comments below the current video is less than 100, all comments will be used. Conversely, negative sampling will be conducted until there are 100 candidates.
@@ 5.1.4. Implementation Details
For the personalized preference alignment stage, we utilize ChatGLM3 (Du et al., 2022; Zeng et al., 2023) as our LLM backbone and perform low-rank adaptation based on LoRA (Hu et al., 2022). We employ the AdamW optimizer with a learning rate of 3e-4 for optimization. With the application of data parallelism and gradient accumulation, the total batch size reaches 128. The loss coefficients, $\lambda$ and $\mu$, are set to 1 and 0.5 respectively. The temperature coefficient $\tau$ is assigned a value of 0.07. The maximum lengths of user video and comment interaction histories are set to 20 for the alignment stage and 50 for the fine-tuning stage. Finally, we conduct a total of 5000 alignment steps. For the recommendation-oriented fine-tuning stage, the embedding dimension is set to 64, and the learning rate is tuned in {0.005, 0.003, 0.001, 0.0005, 0.0001}, which are consistent with all baseline models. The hyperparameter $\eta$ is tuned in {0.1, 0.5 }. More details can be found in our implementation code https://github.com/RUCAIBox/LSVCR.
5.2. Overall Results
Video Recommendation. The overall results of the video recommendation task are shown in Table 2. From the results, we have the following observations: Intuitively incorporating heterogeneous comment interaction behaviors (i.e., DMT) is not better than traditional sequential recommendation models. In contrast, the hypergraph-enhanced Transformer (i.e., MBHT) excels in modeling multi-behavior dependencies and outperforms all traditional baselines. Regarding text-enhanced methods, UniSRec and VQ-Rec achieve better performance, thanks to its contrastive learning pre-training, which leads to high-quality universal sequence representation. As for FDSA and S3-Rec, they do not benefit from the text embeddings. One possible reason is the poor quality of text present on our platform, which requires the adoption of a robust text semantic modeling approach. Finally, LSVCR consistently maintains the best results compared to the baseline methods.
Table 4. Performance comparison of various models on Amazon dataset. “*” denotes that the improvements are significant at the level of 0.01 with paired $t$-test.
Methods Recall@5 Recall@10 NDCG@5 NDCG@10 MRR
Caser 0.4277 0.5554 0.3178 0.3590 0.2985
GRU4Rec 0.5143 0.6368 0.3968 0.4364 0.3743
BERT4Rec 0.4440 0.5665 0.3393 0.3789 0.3210
SASRec 0.5111 0.6276 0.4011 0.4387 0.3802
NARM 0.5121 0.6307 0.3983 0.4367 0.3765
FDSA 0.5261 0.6422 0.4134 0.4509 0.3916
S3Rec 0.5458 0.6631 0.4289 0.4669 0.4060
UniSRect+ID 0.5664 0.6822 0.4480 0.4855 0.4243
VQ-Rec 0.5324 0.6567 0.4267 0.4601 0.3999
LSVCR   w/o Comment 0.6002* 0.7088* 0.4814* 0.5166* 0.4565*
Comment Recommendation. The results of the comment recommendation task are shown in Table 3. From these results, we can find: DSSM, which does not involve user comment interaction history, exhibits inferior performance in comparison to other methods due to the lack of personalized user information. The personalized search methods that apply the title of the current video as a query (i.e., ZAM, TEM) demonstrate better performance than the simple sequential modeling methods (i.e., GRU, ATT, MHA, UniSRecC). Based on the findings, it is evident that the personalized comment recommendation task requires user interests closely aligned with the specific video content, rather than overall generalized user comment preference. By comparing our approach with all baselines, it is clear that LSVCR achieves the most effective results.
Overall, our LSVCR shows significant improvements over all baselines on both video and comment recommendation tasks. This phenomenon indicates that joint modeling of videos and comments can enhance user video and comment preference learning simultaneously. Through personalized preference alignment, we effectively integrate the powerful semantic understanding and knowledge reasoning capabilities of the LLM recommender into the SR model. Furthermore, these results also show that the two-stage training paradigm can coordinate well for improving recommendation.
5.3. Generalizability in Item Recommendation
In this part, we investigate whether the proposed personalized preference alignment is generalizable for other recommender systems. However, to the best of our knowledge, there are no other public datasets consisting of comment data and interactions. Therefore, we employ the “Arts, Crafts and Sewing” subset of Amazon review data (Ni et al., 2019), a general benchmark containing only item titles and user-item interactions, to validate our approach in item recommendation. After processing, the dataset possesses about 20K items and 390K interactions. In this case, the techniques involving comment information in LSVCR will be discarded, including fusion encoders (Section 4.1.3), comment diversification (Section 4.3.1), and video-comment preference contrast (Section 4.3.2). We name this variant of our proposed approach with LSVCR w/o Comment. Additionally, multi-behavior models (i.e., DMT, MBHT) are also not applicable due to the absence of comment interaction behaviors.
The results are shown in Table 4. It can be seen that LSVCR outperforms all baseline models even after discarding the key techniques mentioned above. This indicates that the proposed two-stage training paradigm can effectively integrate the merits of the SR model and the supplemental LLM recommender. Besides, this phenomenon fully demonstrates that our LSVCR can be generalized to more general item recommendation scenarios.
Table 5. Ablation study on two recommendation tasks.